OpenStudy (fanduekisses):

Algebra help( I think it's systems of equations, please help) WILL give TONS of medal! =D

OpenStudy (fanduekisses):

Algebra: Taylor has a total of 43 quarters and nickels in her pocket. if her change is worth $7.75, how many nickels does Taylor have?

OpenStudy (fanduekisses):

C'mon people please help ;'(

OpenStudy (anonymous):

I'll try my best! this is algebra?

OpenStudy (fanduekisses):

Yes :)

OpenStudy (anonymous):

1 or 2?

OpenStudy (fanduekisses):

2

OpenStudy (fanduekisses):

I thin it has something to do with systems of equations (which is in algebra 1 too)

OpenStudy (anonymous):

Okay I'll try since I'm only in Geometry

Directrix (directrix):

Let Q = the number of quarters Taylor has Let N = the number of nickels Taylor has Each quarter has cash value of .25*Q dollars Each nickel has cash value of .05*N dollars So, Q + N = 43 .25Q + .05N = 7.75 Okay, that is the system of equations to solve. @Fanduekisses

OpenStudy (fanduekisses):

Alright so I use the elimination p

OpenStudy (fanduekisses):

Oh that's what I did before but instead of 0.25 i put 25( as a whole, not a cent) lol

OpenStudy (fanduekisses):

Thank you ♥

Directrix (directrix):

Yes, I would also multiply the second equation through by 100 to clear the decimals. Q + N = 43 .25Q + .05N = 7.75 to this: Q + N = 43 25Q + 5N = 775

OpenStudy (fanduekisses):

Then I multiply -5 to the first equation?